# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ Appendix B Tables and Figures
Corn (Grain) CEA 0.199 0.415 0.345 0.336 0.493 0.528 0.498 0.195 0.650 0.547 0.461 0.596 0.565 0.374 0.623 0.446 0.670 0.473 0.704 0.618 0.712 0.673
CWP 0.548 0.391 0.452 0.470 0.560 0.391 0.362 0.455 0.480 0.616 0.442 0.673 0.638 0.635 0.619 0.558 0.534 0.681 0.573 0.620 0.435 0.375
MID 0.228 0.236 0.427 0.215 0.336 0.332 0.490 0.193 0.338 0.442 0.224 0.517 0.616 0.365 0.405 0.482 0.365 0.292 0.482 0.405 0.292 0.156
SEA 0.328 0.497 0.387 0.516 0.476 0.334 0.412 0.563 0.443 0.334 0.207 0.185 0.493 0.366 0.269 0.591 0.796 0.312 0.729 0.759 0.778 0.693
SPL 0.393 0.201 0.214 0.281 0.313 0.271 0.299 0.413 0.310 0.345 0.205 0.294 0.426 0.330 0.343 0.465 0.877 0.426 0.893 0.919 0.844 0.703
SWP 0.612 0.589 0.026 0.558 0.675 0.630 0.602 0.514 0.453 0.212 0.750 0.621 0.580 0.653 0.487 0.653 0.620 0.630 0.756 0.741 0.689 0.497
Wheat (Winter) CEA 0.468 0.589 0.418 0.608 0.308 0.502 0.496 0.349 0.539 0.465 0.338 0.429 0.634 0.455 0.681 0.583 0.639 0.559 0.705 0.724 0.767 0.677
CWP 0.475 0.753 0.216 0.621 0.327 0.390 0.478 0.595 0.492 0.635 0.453 0.811 0.700 0.698 0.646 0.559 0.541 0.496 0.498 0.533 0.392 0.349
MID 0.372 0.411 0.622 0.368 0.351 0.357 0.281 0.279 0.532 0.281 0.252 0.311 0.296 0.364 0.461 0.432 0.461 0.328 0.364 0.432 0.437 0.328
SEA 0.506 0.423 0.354 0.667 0.287 0.274 0.415 0.451 0.361 0.184 0.360 0.527 0.354 0.272 0.328 0.189 0.667 0.393 0.832 0.652 0.669 0.588
SPL 0.386 0.453 0.267 0.355 0.579 0.390 0.463 0.396 0.366 0.428 0.315 0.341 0.608 0.375 0.280 0.423 0.800 0.325 0.792 0.757 0.570 0.734
SWP 0.429 0.459 0.284 0.540 0.382 0.485 0.463 0.317 0.325 0.389 0.438 0.536 0.584 0.585 0.698 0.712 0.487 0.530 0.525 0.463 0.390 0.532
Soybeans CEA 0.201 0.427 0.387 0.423 0.595 0.644 0.579 0.167 0.608 0.559 0.447 0.643 0.597 0.430 0.545 0.534 0.718 0.453 0.602 0.682 0.652 0.573
MID 0.323 0.375 0.494 0.250 0.397 0.396 0.548 0.406 0.672 0.360 0.277 0.593 0.620 0.584 0.435 0.616 0.616 0.344 0.435 0.584 0.344 0.423
SEA 0.358 0.375 0.469 0.465 0.562 0.395 0.488 0.383 0.364 0.443 0.135 0.193 0.559 0.479 0.383 0.728 0.723 0.309 0.685 0.729 0.659 0.472
SPL 0.426 0.223 0.334 0.272 0.408 0.268 0.520 0.249 0.414 0.362 0.413 0.440 0.507 0.423 0.400 0.580 0.826 0.582 0.842 0.786 0.768 0.642
@@
Table B5: MAPE values of GAMs for each crop and region (Test Set)
Crop Region Model
@@
Corn (Grain) CEA 0.228 0.259 0.288 0.289 0.214 0.210 0.185 0.302 0.263 0.303 0.259 0.313 0.255 0.274 0.201 0.289 0.075 0.163 0.102 0.131 0.224 0.108
CWP 0.294 0.289 0.313 0.400 0.562 0.314 0.243 0.402 0.230 0.305 0.290 0.256 0.261 0.246 0.215 0.288 0.241 0.165 0.172 0.243 0.190 0.302
MID 0.281 0.329 0.332 0.314 0.251 0.304 0.295 0.359 0.339 0.332 0.389 0.310 0.307 0.305 0.234 0.228 0.305 0.279 0.234 0.228 0.347 0.279
SEA 0.138 0.248 0.459 0.214 0.254 0.335 0.226 0.295 0.272 0.364 0.330 0.262 0.214 0.228 0.201 0.450 0.323 0.280 0.272 0.183 0.194 0.175
SPL 0.129 0.230 0.275 0.238 0.156 0.264 0.184 0.170 0.222 0.259 0.171 0.164 0.178 0.196 0.231 0.172 0.085 0.226 0.079 0.167 0.118 0.151
SWP 0.138 0.104 0.214 0.105 0.159 0.178 0.068 0.097 0.103 0.221 0.260 0.228 0.139 0.155 0.069 0.118 0.072 0.130 0.114 0.228 0.129 0.368
Wheat (Winter) CEA 0.172 0.197 0.354 0.176 0.200 0.205 0.161 0.235 0.195 0.239 0.235 0.210 0.240 0.239 0.150 0.269 0.042 0.165 0.081 0.088 0.040 0.148
CWP 0.456 0.207 0.220 0.187 0.168 0.225 0.175 0.165 0.257 0.148 0.240 0.443 0.160 0.117 0.090 0.163 0.168 0.186 0.134 0.170 0.198 0.160
MID 0.188 0.294 0.293 0.288 0.226 0.265 0.248 0.382 0.286 0.265 0.341 0.320 0.302 0.311 0.281 0.340 0.311 0.282 0.281 0.340 0.282 0.363
SEA 0.229 0.222 0.399 0.380 0.271 0.275 0.262 0.319 0.204 0.290 0.228 0.305 0.287 0.229 0.150 0.307 0.104 0.225 0.074 0.085 0.066 0.132
SPL 0.171 0.238 0.217 0.212 0.139 0.199 0.174 0.145 0.175 0.228 0.168 0.184 0.181 0.192 0.208 0.201 0.131 0.215 0.109 0.111 0.151 0.104
SWP 0.134 0.110 0.109 0.112 0.073 0.069 0.079 0.167 0.086 0.144 0.094 0.125 0.039 0.145 0.067 0.092 0.153 0.044 0.068 0.103 0.161 0.217
Soybeans CEA 0.193 0.205 0.251 0.235 0.138 0.159 0.140 0.271 0.244 0.257 0.206 0.262 0.205 0.225 0.175 0.224 0.099 0.141 0.097 0.089 0.123 0.155
MID 0.234 0.290 0.378 0.273 0.214 0.268 0.231 0.314 0.288 0.282 0.345 0.286 0.258 0.278 0.248 0.233 0.233 0.243 0.248 0.278 0.243 0.379
SEA 0.193 0.279 0.339 0.262 0.221 0.276 0.234 0.294 0.299 0.305 0.324 0.263 0.235 0.240 0.186 0.374 0.093 0.248 0.106 0.107 0.108 0.238
SPL 0.173 0.221 0.265 0.221 0.211 0.249 0.185 0.169 0.247 0.236 0.230 0.145 0.203 0.184 0.194 0.182 0.077 0.199 0.104 0.108 0.141 0.160
[…]
Crop Region Model

### v2

@@ Appendix B Tables
Train Corn CEA 0.047 0.053 0.044 0.051 0.042 0.047 0.043 0.038 0.046 0.048 0.042 0.039 0.055 0.043 0.047 0.039 0.048 0.044 0.044 0.040 0.041 0.053
CWP 0.039 0.035 0.029 0.036 0.037 0.034 0.039 0.035 0.037 0.025 0.034 0.032 0.029 0.032 0.025 0.029 0.030 0.031 0.027 0.035 0.034 0.032
MID 0.028 0.034 0.037 0.025 0.032 0.034 0.035 0.036 0.037 0.044 0.032 0.044 0.040 0.044 0.042 0.044 0.044 0.053 0.042 0.044 0.053 0.037
SEA 0.037 0.057 0.054 0.035 0.036 0.047 0.040 0.040 0.058 0.055 0.043 0.046 0.046 0.049 0.053 0.053 0.043 0.051 0.056 0.047 0.043 0.055
SPL 0.137 0.146 0.160 0.139 0.147 0.135 0.132 0.166 0.130 0.127 0.157 0.127 0.130 0.124 0.116 0.118 0.062 0.137 0.058 0.061 0.066 0.116
SWP 0.137 0.154 0.151 0.162 0.133 0.090 0.123 0.134 0.155 0.145 0.126 0.110 0.098 0.088 0.139 0.088 0.095 0.090 0.084 0.075 0.089 0.132
Wheat CEA 0.038 0.035 0.032 0.035 0.035 0.029 0.032 0.040 0.031 0.038 0.039 0.035 0.034 0.029 0.035 0.033 0.031 0.041 0.037 0.033 0.036 0.037
CWP 0.057 0.052 0.047 0.058 0.051 0.049 0.053 0.050 0.052 0.040 0.052 0.033 0.036 0.039 0.042 0.038 0.042 0.039 0.043 0.040 0.047 0.044
MID 0.043 0.041 0.036 0.046 0.042 0.036 0.037 0.042 0.045 0.035 0.036 0.039 0.040 0.040 0.042 0.038 0.040 0.042 0.042 0.038 0.042 0.041
SEA 0.044 0.047 0.045 0.039 0.038 0.031 0.032 0.047 0.048 0.051 0.051 0.043 0.046 0.041 0.045 0.039 0.033 0.039 0.044 0.039 0.043 0.044
SPL 0.041 0.045 0.040 0.041 0.037 0.040 0.045 0.036 0.050 0.039 0.044 0.042 0.043 0.040 0.043 0.045 0.042 0.042 0.042 0.047 0.045 0.050
SWP 0.063 0.061 0.054 0.069 0.063 0.053 0.044 0.064 0.061 0.055 0.068 0.041 0.043 0.051 0.050 0.038 0.051 0.042 0.049 0.044 0.048 0.057
Soybeans CEA 0.052 0.059 0.037 0.061 0.037 0.044 0.038 0.046 0.061 0.049 0.050 0.044 0.057 0.047 0.051 0.046 0.059 0.047 0.053 0.042 0.047 0.059
MID 0.029 0.035 0.029 0.026 0.027 0.027 0.029 0.034 0.028 0.033 0.032 0.031 0.032 0.032 0.029 0.033 0.032 0.035 0.029 0.033 0.035 0.035
SEA 0.038 0.057 0.051 0.040 0.040 0.043 0.049 0.037 0.045 0.047 0.042 0.050 0.058 0.051 0.050 0.056 0.042 0.051 0.048 0.043 0.048 0.055
SPL 0.036 0.045 0.037 0.044 0.036 0.040 0.038 0.046 0.046 0.044 0.047 0.040 0.036 0.046 0.037 0.043 0.039 0.047 0.041 0.043 0.046 0.049
Test Corn CEA 0.109 0.129 0.101 0.106 0.102 0.103 0.108 0.103 0.112 0.100 0.100 0.105 0.119 0.111 0.113 0.108 0.112 0.115 0.119 0.105 0.098 0.129
CWP 0.043 0.049 0.038 0.044 0.041 0.042 0.042 0.037 0.042 0.038 0.042 0.037 0.037 0.039 0.040 0.043 0.039 0.049 0.045 0.043 0.041 0.046
MID 0.055 0.053 0.047 0.052 0.053 0.055 0.050 0.049 0.049 0.056 0.047 0.050 0.051 0.056 0.060 0.060 0.060 0.058 0.056 0.060 0.058 0.064
SEA 0.074 0.074 0.071 0.072 0.070 0.068 0.078 0.073 0.086 0.067 0.076 0.073 0.066 0.076 0.080 0.087 0.082 0.087 0.076 0.084 0.081 0.093
SPL 0.235 0.273 0.293 0.239 0.256 0.268 0.230 0.277 0.271 0.292 0.247 0.258 0.246 0.230 0.258 0.242 0.099 0.258 0.082 0.097 0.119 0.152
SWP 0.163 0.186 0.306 0.129 0.162 0.174 0.157 0.144 0.184 0.285 0.124 0.135 0.157 0.114 0.155 0.178 0.112 0.176 0.119 0.148 0.188 0.179
Wheat CEA 0.059 0.060 0.067 0.065 0.068 0.068 0.070 0.069 0.053 0.067 0.066 0.060 0.062 0.071 0.075 0.059 0.053 0.079 0.053 0.079 0.074 0.065
CWP 0.102 0.090 0.103 0.115 0.107 0.085 0.093 0.081 0.094 0.094 0.089 0.095 0.095 0.086 0.094 0.092 0.100 0.086 0.092 0.105 0.086 0.099
MID 0.076 0.073 0.073 0.067 0.066 0.071 0.062 0.075 0.076 0.065 0.063 0.072 0.068 0.070 0.070 0.072 0.070 0.090 0.070 0.072 0.090 0.078
SEA 0.072 0.088 0.056 0.083 0.075 0.073 0.077 0.085 0.079 0.067 0.072 0.065 0.060 0.080 0.083 0.069 0.072 0.076 0.070 0.078 0.084 0.075
SPL 0.094 0.089 0.083 0.083 0.084 0.081 0.073 0.087 0.072 0.073 0.086 0.092 0.077 0.087 0.081 0.078 0.080 0.077 0.062 0.088 0.083 0.096
SWP 0.154 0.151 0.139 0.154 0.159 0.146 0.148 0.156 0.149 0.136 0.147 0.160 0.139 0.141 0.150 0.157 0.148 0.146 0.144 0.157 0.154 0.172
Soybeans CEA 0.105 0.116 0.085 0.107 0.083 0.094 0.103 0.098 0.116 0.107 0.097 0.102 0.106 0.104 0.102 0.112 0.086 0.097 0.101 0.101 0.098 0.114
MID 0.061 0.058 0.047 0.064 0.046 0.047 0.053 0.056 0.057 0.052 0.059 0.045 0.059 0.053 0.052 0.052 0.046 0.059 0.052 0.052 0.059 0.056
SEA 0.103 0.133 0.101 0.123 0.102 0.101 0.093 0.112 0.109 0.099 0.120 0.106 0.117 0.102 0.110 0.104 0.096 0.092 0.109 0.100 0.111 0.128
SPL 0.077 0.094 0.076 0.078 0.076 0.085 0.082 0.070 0.086 0.070 0.077 0.076 0.075 0.067 0.059 0.063 0.070 0.071 0.088 0.067 0.071 0.084
@@
Table B5: MAPE values of LGBM for each crop, model, and region
Set Crop Region Model
@@
Train Corn CEA 0.093 0.100 0.088 0.095 0.083 0.098 0.087 0.086 0.095 0.094 0.094 0.089 0.093 0.091 0.092 0.093 0.086 0.094 0.086 0.094 0.094 0.100
CWP 0.067 0.064 0.064 0.064 0.063 0.063 0.065 0.063 0.065 0.060 0.065 0.058 0.062 0.062 0.059 0.063 0.061 0.065 0.059 0.065 0.065 0.060
MID 0.076 0.080 0.085 0.073 0.078 0.077 0.081 0.076 0.078 0.081 0.077 0.083 0.085 0.085 0.083 0.080 0.085 0.087 0.083 0.080 0.087 0.082
SEA 0.105 0.111 0.111 0.097 0.103 0.107 0.109 0.104 0.113 0.116 0.107 0.102 0.101 0.109 0.110 0.107 0.103 0.105 0.109 0.102 0.097 0.116
SPL 0.280 0.286 0.295 0.285 0.296 0.285 0.273 0.287 0.280 0.271 0.286 0.274 0.269 0.262 0.264 0.274 0.208 0.268 0.206 0.220 0.214 0.249
SWP 0.302 0.326 0.352 0.307 0.319 0.297 0.296 0.304 0.336 0.345 0.304 0.284 0.274 0.292 0.293 0.275 0.237 0.299 0.283 0.262 0.302 0.302
Wheat CEA 0.071 0.070 0.066 0.071 0.063 0.062 0.063 0.069 0.071 0.066 0.068 0.066 0.068 0.064 0.066 0.067 0.066 0.069 0.068 0.071 0.069 0.070
CWP 0.106 0.097 0.098 0.103 0.103 0.094 0.100 0.096 0.101 0.081 0.095 0.090 0.082 0.093 0.089 0.091 0.092 0.091 0.090 0.092 0.100 0.097
MID 0.078 0.078 0.072 0.077 0.076 0.072 0.072 0.079 0.078 0.074 0.078 0.077 0.074 0.075 0.079 0.071 0.075 0.074 0.079 0.071 0.074 0.076
SEA 0.087 0.092 0.084 0.090 0.085 0.083 0.080 0.092 0.090 0.093 0.091 0.083 0.088 0.086 0.088 0.086 0.089 0.083 0.091 0.084 0.085 0.092
SPL 0.075 0.077 0.075 0.074 0.073 0.072 0.075 0.078 0.078 0.075 0.076 0.074 0.076 0.074 0.075 0.073 0.074 0.076 0.074 0.076 0.075 0.078
SWP 0.118 0.120 0.122 0.116 0.118 0.116 0.115 0.121 0.119 0.119 0.119 0.113 0.112 0.115 0.112 0.113 0.108 0.100 0.109 0.108 0.110 0.120
Soybeans CEA 0.101 0.104 0.091 0.101 0.092 0.094 0.091 0.094 0.101 0.095 0.097 0.092 0.096 0.096 0.097 0.100 0.098 0.098 0.098 0.094 0.097 0.104
MID 0.057 0.062 0.055 0.060 0.056 0.055 0.058 0.058 0.059 0.058 0.059 0.058 0.058 0.061 0.057 0.060 0.061 0.059 0.057 0.060 0.059 0.060
SEA 0.100 0.112 0.115 0.102 0.106 0.110 0.110 0.103 0.106 0.120 0.105 0.107 0.116 0.114 0.116 0.115 0.102 0.111 0.110 0.106 0.109 0.115
SPL 0.079 0.082 0.077 0.081 0.074 0.084 0.079 0.083 0.083 0.083 0.081 0.080 0.075 0.081 0.082 0.082 0.083 0.085 0.082 0.086 0.084 0.083
Test Corn CEA 0.111 0.112 0.100 0.117 0.101 0.105 0.102 0.115 0.107 0.101 0.110 0.100 0.109 0.116 0.112 0.114 0.113 0.114 0.103 0.105 0.110 0.110
CWP 0.035 0.034 0.037 0.037 0.037 0.036 0.036 0.035 0.036 0.040 0.036 0.039 0.038 0.036 0.036 0.035 0.039 0.037 0.038 0.036 0.034 0.038
MID 0.055 0.052 0.051 0.048 0.047 0.047 0.056 0.052 0.054 0.052 0.055 0.049 0.055 0.054 0.053 0.055 0.053 0.053 0.054 0.055 0.053 0.054
SEA 0.071 0.079 0.069 0.079 0.072 0.068 0.075 0.079 0.081 0.077 0.078 0.079 0.073 0.078 0.080 0.074 0.079 0.073 0.073 0.072 0.074 0.087
SPL 0.270 0.290 0.293 0.274 0.278 0.287 0.261 0.266 0.255 0.287 0.273 0.267 0.264 0.261 0.282 0.274 0.187 0.282 0.187 0.195 0.206 0.220
SWP 0.198 0.226 0.298 0.200 0.230 0.249 0.226 0.194 0.234 0.287 0.187 0.227 0.263 0.210 0.235 0.238 0.203 0.230 0.164 0.214 0.230 0.214
Wheat CEA 0.067 0.063 0.066 0.061 0.065 0.066 0.066 0.065 0.066 0.064 0.063 0.064 0.062 0.059 0.063 0.061 0.061 0.064 0.065 0.062 0.064 0.066
CWP 0.090 0.093 0.099 0.101 0.093 0.094 0.095 0.097 0.102 0.106 0.098 0.099 0.100 0.112 0.106 0.103 0.105 0.101 0.090 0.100 0.100 0.103
MID 0.066 0.069 0.066 0.067 0.072 0.066 0.061 0.067 0.064 0.068 0.067 0.069 0.069 0.066 0.066 0.065 0.066 0.072 0.069 0.065 0.071 0.072
SEA 0.065 0.067 0.067 0.062 0.061 0.065 0.064 0.065 0.068 0.064 0.065 0.073 0.063 0.060 0.065 0.070 0.062 0.067 0.059 0.068 0.068 0.068
SPL 0.079 0.077 0.075 0.073 0.074 0.070 0.071 0.077 0.069 0.070 0.076 0.070 0.074 0.066 0.073 0.072 0.075 0.075 0.074 0.075 0.078 0.079
SWP 0.164 0.148 0.160 0.157 0.159 0.151 0.153 0.159 0.152 0.157 0.151 0.157 0.142 0.151 0.151 0.151 0.167 0.149 0.146 0.168 0.164 0.176
Soybeans CEA 0.101 0.102 0.100 0.103 0.090 0.099 0.091 0.103 0.101 0.093 0.104 0.101 0.094 0.101 0.105 0.104 0.091 0.107 0.106 0.101 0.105 0.097
MID 0.055 0.053 0.047 0.053 0.052 0.051 0.054 0.053 0.053 0.051 0.054 0.050 0.049 0.051 0.050 0.054 0.042 0.051 0.050 0.054 0.051 0.052
SEA 0.116 0.129 0.109 0.129 0.114 0.123 0.119 0.129 0.136 0.114 0.129 0.116 0.126 0.129 0.119 0.123 0.110 0.106 0.125 0.131 0.126 0.137
SPL 0.079 0.081 0.076 0.077 0.079 0.073 0.076 0.080 0.083 0.073 0.077 0.079 0.083 0.075 0.080 0.073 0.081 0.081 0.078 0.078 0.073 0.079
Table B6: MAPE values of XGB-C for each crop, model, and region
Set Crop Region Model